OpenStudy (anonymous):

what is the sum 5y6+9y6

OpenStudy (texaschic101):

14y^6

OpenStudy (anonymous):

how did you get that answer ?

OpenStudy (texaschic101):

what you have is like terms...when adding like terms, just add the coefficients. Examples : 2n + 3n = 5n 3x^2 + 2x^2 = 5x^2 2x^3 + 3x^3 = 5x^3 so for 5y^6 + 9y^6 ....add 5 and 9 and put it with y^6 and you get 14y^6 understand ?

OpenStudy (texaschic101):

now this is only for adding/subtracting. Not for multiplying/dividing

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

anytime :)